# Improving Genetic Programming for Symbolic Regression with Equality Graphs

### v1

@@ 6. Results and Discussion
6.1. Hyperparameter tuning
Applying the Wilcoxon rank test to the obtained results using , we observed that only tinyGP expressed a significant difference in their results (-value ) favoring generations and population size . There was no significant difference observed in any algorithm, thus, we chose the configuration which found the best result for most datasets. With this criteria, PySR and Operon was set to generations and eggp with generations.
6.2. New SRBench datasets
@@
In Fig. 3 we can see the performance plots for each SRBench dataset of the best solution according to the highest on the training set. The x-axis of these plots represents the measured on the test set, and the y-axis shows the percentage of runs that the algorithm found an equal or larger than . The ideal algorithm would cover the whole area from to . As we can see from these plots, eggp always covers an area similar or better than the competing algorithms. Considering the area under the curve values (AUC), eggp covered the most area in out of the datasets, Operon in , PySR in and tinyGP in . It is also evident from these plots how eggp displays a significant improvement over tinyGP.
@@
In Table 3 we can see the ranks when considering the median of the test set and the AUC. We can see that, using this criteria, eggp is consistently ranked second, while Operon and PySR take turns in the first and third place. The statistical test reveals that we can reject the null hypothesis when comparing to PySR with the alternative of having greater median rank. On the other hand, for the AUC values, we can see that eggp is greater than the other algorithms on average while rejecting the null hypotheses for each comparison. Unlike the median of the , the AUC is the average weighted by the probability of obtaining that value or greater, acting as a reliability measure of obtaining that value or greater.
| dataset | eggp | Operon | PySR | tinyGP |
| 192_vineyard | 1 | 4 | 3 | 2 |
| 210_cloud | 1 | 4 | 2 | 3 |
| 522_pm10 | 3 | 4 | 1 | 2 |
| 557_analcatdata_apnea1 | 2 | 4 | 3 | 1 |
| 579_fri_c0_250_5 | 2 | 1 | 3 | 4 |
| 606_fri_c2_1000_10 | 1 | 2 | 3 | 4 |
| 650_fri_c0_500_50 | 2 | 1 | 3 | 4 |
| 678_visualizing_environmental | 1 | 4 | 2 | 3 |
| 1028_SWD | 4 | 1 | 2 | 3 |
| 1089_USCrime | 1 | 3 | 4 | 2 |
| 1193_BNG_lowbwt | 2 | 1 | 3 | 4 |
| 1199_BNG_echoMonths | 4 | 3 | 1 | 2 |
| mean | 2.00 | 2.67 | 2.50 | 2.83 |
| -value | ||||
| 192_vineyard | 0.24 | 0.08 | 0.16 | 0.17 |
| 210_cloud | 0.71 | 0.34 | 0.71 | 0.58 |
| 522_pm10 | 0.16 | 0.14 | 0.19 | 0.18 |
| 557_analcatdata_apnea1 | 0.75 | 0.40 | 0.58 | 0.68 |
| 579_fri_c0_250_5 | 0.91 | 0.95 | 0.83 | 0.79 |
| 606_fri_c2_1000_10 | 0.96 | 0.96 | 0.82 | 0.70 |
| 650_fri_c0_500_50 | 0.87 | 0.92 | 0.83 | 0.40 |
| 678_visualizing_environmental | 0.28 | 0.08 | 0.27 | 0.20 |
| 1028_SWD | 0.38 | 0.39 | 0.39 | 0.36 |
| 1089_USCrime | 0.68 | 0.59 | 0.59 | 0.59 |
| 1193_BNG_lowbwt | 0.56 | 0.57 | 0.56 | 0.55 |
| 1199_BNG_echoMonths | 0.40 | 0.30 | 0.43 | 0.35 |
| mean | 0.58 | 0.48 | 0.53 | 0.46 |
| -value |
6.3. Real-world datasets
@@
In Fig. 4 we observe a similar behavior with eggp covering an area close to or better than the best competing algorithm. Considering AUC, it found out of best results, while PySR only , Operon and tinyGP . Again, considering the ranks on the median of the (Table 4), it consistently achieved second place, but in this set, Operon was ranked first for most of the datasets. We should notice, though, that eggp results were always close to the best competing algorithm, while even Operon misbehaved in two datasets (flow and niku-2). When observing the statistical test results, we can conclude that there are no significant differences between the eggp and Operon, but the hypotheses of eggp being equivalent to PySR and tinyGP can be rejected.
| dataset | eggp | Operon | PySR | tinyGP |
| Chemical_1_tower | 2 | 1 | 4 | 3 |
| Chemical_2_competition | 2 | 1 | 4 | 3 |
| Flow_stress_phip0.1 | 1 | 4 | 2 | 3 |
| Friction_dyn_one-hot | 2 | 1 | 4 | 3 |
| Friction_stat_one-hot | 2 | 1 | 4 | 3 |
| Nasa_battery_1_10min | 2 | 1 | 4 | 3 |
| Nasa_battery_2_20min | 2 | 3 | 4 | 1 |
| Nikuradse_1 | 2 | 1 | 4 | 3 |
| Nikuradse_2 | 3 | 1 | 4 | 2 |
| mean | 2.00 | 1.56 | 3.78 | 2.67 |
| -value | ||||
| Chemical_1_tower | 0.86 | 0.89 | 0.67 | 0.81 |
| Chemical_2_competition | 0.58 | 0.66 | 0.07 | 0.37 |
| Flow_stress_phip0.1 | 0.99 | 0.51 | 0.98 | 0.88 |
| Friction_dyn_one-hot | 0.81 | 0.85 | 0.66 | 0.74 |
| Friction_stat_one-hot | 0.77 | 0.82 | 0.30 | 0.65 |
| Nasa_battery_1_10min | 0.96 | 0.96 | 0.85 | 0.93 |
| Nasa_battery_2_20min | 0.97 | 0.95 | 0.97 | 0.98 |
| Nikuradse_1 | 0.85 | 0.89 | 0.62 | 0.75 |
| Nikuradse_2 | 0.98 | 0.83 | 0.95 | 0.95 |
| mean | 0.87 | 0.82 | 0.67 | 0.79 |
| -value |
6.4. Hypervolume
Table 5 shows the average and standard deviation of the hypervolume of the final Pareto fronts. For SRBench, eggp covered a similar area to the other algorithms. PySR returned the largest hypervolume in most of the SRBench datasets. For the real-world datasets, eggp covered the most hypervolume overall.
| dataset | egraphGP | operon | PySR |
| srbench | |||
| 192 | |||
| 210 | |||
| 522 | |||
| 557 | |||
| 579 | |||
| 606 | |||
| 650 | |||
| 678 | |||
| 1028 | |||
| 1089 | |||
| 1193 | |||
| 1199 | |||
| real-world | |||
| Chemical_1 | |||
| Chemical_2 | |||
| Flow_stress | |||
| Friction_dyn | |||
| Friction_stat | |||
| Nasa_1 | |||
| Nasa_2 | |||
| Nikuradse_1 | |||
| Nikuradse_2 | |||
6.5. Computational time
@@
Regarding the runtime, since Operon is the fastest symbolic regression implementation, as noted in (Burlacu et al., 2020b), we calculated the ratio between the average runtime of each algorithm to Operon. Fig. 5 shows the relative runtime per dataset. In this plot we can see that eggp and tinyGP were both between to times slower than Operon. PySR varied from to times the runtime of Operon depending on the dataset. The higher ratios were observed on high-dimensional or larger datasets. We should stress that all algorithms were constrained to run with a single thread, thus both Operon and PySR runtime could be smaller when exploiting multi-threading.
@@ 7. Conclusions
In this paper we explored the use of e-graphs and equality saturation as a mechanism to keep track of the history of the search engine of symbolic regression and exploiting its pattern matching capabilities to propose a variation to the traditional subtree crossover and mutation that increases the probability of generating novel expressions.
The e-graph data structure efficiently stores and queries for parts of expressions. Exploiting this capability, we modified the subtree operators to only sample subtrees that would form unvisited expressions. The expectation is that this simple modification would render a significant improvement in the search procedure.
We tested the proposed algorithm, called eggp, in different benchmarks from the literature and compared with a simple GP with the original subtree operators, and two state-of-the-art algorithms, PySR and Operon. The results showed that the modified operators are capable of improving the performance of a simple GP to compete with the state-of-the-art. The main highlight of this approach is that it consistently performs equal or better than the best competing approach.
The main limitations of these experiments lie in the use of default or reasonable values for the hyperparameters. We should notice that eggp contains eight parameters that should be adjusted accordingly in a practical scenario. Meanwhile, PySR contains about hyperparameters that may affect the algorithm performance888we are not considering hyperparameters that expect prior knowledge. and Operon contains about hyperparameters. With careful experimentation, both PySR and Operon could possibly achieve similar results. Regarding the runtime, eggp is consistently faster than PySR but significantly slower than Operon. When comparing with tinyGP, we can see that the use of e-graph and equality saturation does not increase the runtime significantly.
In conclusion, the expressiveness and capabilities of the e-graph data structure enabled us to make a simple modification to the original subtree operators while significantly improving the performance of GP for symbolic regression. This allowed us to obtain a more robust algorithm delivering better performance more reliably than state-of-the-art implementations. As for the next steps, this same modifications can be applied to any other mutation operator used by the state-of-the-art algorithms. Not only that, but the e-graph opens up many new possibilities for improving the search as it allows us to query expressions with a combination of properties, which can translate to diversity-preserving population and easy to integrate prior knowledge (Kronberger et al., 2022; Reuter et al., 2024). In addition, the storage of the search history allows us to analyze the learned building blocks and exploit this information to generate new solutions. 999The authors proudly made no use of LLMs for this work.
@@ Acknowledgements.
F.O.F. is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) grant 301596/2022-0. G.K. is supported by the Austrian Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation and Technology, the Federal Ministry for Labour and Economy, and the regional government of Upper Austria within the COMET project ProMetHeus (904919) supported by the Austrian Research Promotion Agency (FFG). No LLMs were used for this work.References
@@
- Bartlett et al. (2023) Deaglan J Bartlett, Harry Desmond, and Pedro G Ferreira. 2023. Exhaustive symbolic regression. IEEE Transactions on Evolutionary Computation (2023).
- Burlacu et al. (2019) Bogdan Burlacu, Michael Affenzeller, Gabriel Kronberger, and Michael Kommenda. 2019. Online diversity control in symbolic regression via a fast hash-based tree similarity measure. In 2019 IEEE congress on evolutionary computation (CEC). IEEE, 2175–2182.
- Burlacu et al. (2020a) Bogdan Burlacu, Lukas Kammerer, Michael Affenzeller, and Gabriel Kronberger. 2020a. Hash-based tree similarity and simplification in genetic programming for symbolic regression. In Computer Aided Systems Theory–EUROCAST 2019: 17th International Conference, Las Palmas de Gran Canaria, Spain, February 17–22, 2019, Revised Selected Papers, Part I 17. Springer, 361–369.
- Burlacu et al. (2020b) Bogdan Burlacu, Gabriel Kronberger, and Michael Kommenda. 2020b. Operon C++ an efficient genetic programming framework for symbolic regression. In Proceedings of the 2020 Genetic and Evolutionary Computation Conference Companion. 1562–1570.
- Cao et al. ([n. d.]) Haotian Cao, Garrett W Merz, Kyle Cranmer, and Gary Shiu. [n. d.]. Learning Conformal Field Theory with Symbolic Regression: Recovering the Symbolic Expressions for the Energy Spectrum. ([n. d.]).
@@
- de Franca et al. (2024) Fabricio O de Franca, Marco Virgolin, M Kommenda, MS Majumder, M Cranmer, G Espada, L Ingelse, A Fonseca, M Landajuela, B Petersen, et al. 2024. SRBench++: Principled benchmarking of symbolic regression with domain-expert interpretation. IEEE transactions on evolutionary computation (2024).
- Deb et al. (2000) Kalyanmoy Deb, Samir Agrawal, Amrit Pratap, and Tanaka Meyarivan. 2000. A fast elitist non-dominated sorting genetic algorithm for multi-objective optimization: NSGA-II. In Parallel Problem Solving from Nature PPSN VI: 6th International Conference Paris, France, September 18–20, 2000 Proceedings 6. Springer, 849–858.
- Ebner et al. (2001) Marc Ebner, Mark Shackleton, and Rob Shipman. 2001. How neutral networks influence evolvability. Complex. 7, 2 (Nov. 2001), 19–33. https://doi.org/10.1002/cplx.10021
- Gustafson et al. (2005) Steven Gustafson, Edmund K Burke, and Natalio Krasnogor. 2005. On improving genetic programming for symbolic regression. In 2005 IEEE Congress on Evolutionary Computation, Vol. 1. IEEE, 912–919.
- Hu and Banzhaf (2018) Ting Hu and Wolfgang Banzhaf. 2018. Neutrality, robustness, and evolvability in genetic programming. Genetic Programming Theory and Practice XIV (2018), 101–117.
- Hu et al. (2023) Ting Hu, Gabriela Ochoa, and Wolfgang Banzhaf. 2023. Phenotype search trajectory networks for linear genetic programming. In European Conference on Genetic Programming (Part of EvoStar). Springer, 52–67.
- Keller and Banzhaf (1999) Robert E Keller and Wolfgang Banzhaf. 1999. The evolution of genetic code in genetic programming. In Proceedings of the Genetic and Evolutionary Computation Conference, Vol. 2. Morgan Kaufmann Orlando, 1077–1082.
@@
- Kronberger and de França (2024) Gabriel Kronberger and Fabrício Olivetti de França. 2024. Effects of reducing redundant parameters in parameter optimization for symbolic regression using genetic programming. Journal of Symbolic Computation (2024), 102413.
- Kronberger et al. (2022) Gabriel Kronberger, Fabricio Olivetti de França, Bogdan Burlacu, Christian Haider, and Michael Kommenda. 2022. Shape-constrained symbolic regression—improving extrapolation with prior knowledge. Evolutionary Computation 30, 1 (2022), 75–98.
@@
- Lamb and Harrington ([n. d.]) Kara D Lamb and Jerry Y Harrington. [n. d.]. Discovering How Ice Crystals Grow Using Neural ODE’s and Symbolic Regression. ([n. d.]).
@@
- Seidyo Imai Aldeia et al. (2024) Guilherme Seidyo Imai Aldeia, Fabricio Olivetti de Franca, and William G La Cava. 2024. Inexact Simplification of Symbolic Regression Expressions with Locality-sensitive Hashing. arXiv e-prints (2024), arXiv–2404.
@@
- Virgolin and Pissis ([n. d.]) Marco Virgolin and Solon P Pissis. [n. d.]. Symbolic Regression is NP-hard. Transactions on Machine Learning Research ([n. d.]).

### v2

@@ 6. Results and Discussion
In Fig. 3 we can see the performance plots for the SRBench datasets considering the best solution of each run according to the highest on the training set. The x-axis of these plots represents the measured on the test set, and the y-axis shows the percentage of runs that the algorithm found an equal or larger than . The ideal algorithm would cover the whole area from to . From these plots we can see that in at least datasets (, and ) every algorithm reliably achieve the same . In other datasets (, , ) both versions of eggp maintain this reliability (i.e., achieves the same score in almost every execution) while the other algorithms either achieve a lower score or fails in some execution. The failing executions can be identified as , for dataset every algorithm fails between to of the times.
@@
In Table 3 we can see the ranks when considering the median of the test set and the AUC. Using this criteria, eggpmo have an average rank of , while Operon comes next ranked on average. Considering eggpso and tinyGP, there is a slightly decrease in the average rank when using the proposed operators. The statistical test reveals that we can reject the null hypothesis when comparing to PySR with the alternative of having greater median rank. On the other hand, for the AUC values, we can see that eggpmo is greater than the other algorithms on average while rejecting the null hypotheses for each comparison, except Operon. Unlike the median of the , the AUC is the average weighted by the probability of obtaining that value or greater, acting as a reliability measure. Also in this table, we can see that, on average, eggp (both versions) consistently return smaller models than the competing algorithms. There are two possible reasons for this behavior: i) as we apply equality saturation after inserting each expression into the e-graph, they can result in a simplified version of the inserted expression, ii) during the insertion, it automatically eliminates some of the redundant parameters, avoiding the issue reported in(de Franca and Kronberger, 2023).
| dataset | eggpmo | eggpso | Operon | PySR | tinyGP |
| 192 | |||||
| 210 | |||||
| 522 | |||||
| 557 | |||||
| 579 | |||||
| 606 | |||||
| 650 | |||||
| 678 | |||||
| 1028 | |||||
| 1089 | |||||
| 1193 | |||||
| 1199 | |||||
| mean | |||||
| -value | 0.05 | 0.21 | 0.02 | 0.11 | |
| 192 | |||||
| 210 | |||||
| 522 | |||||
| 557 | |||||
| 579 | |||||
| 606 | |||||
| 650 | |||||
| 678 | |||||
| 1028 | |||||
| 1089 | |||||
| 1193 | |||||
| 1199 | |||||
| mean | |||||
| -value | 0.05 | 0.23 | 0.05 | 0.01 | |
| 192 | |||||
| 210 | |||||
| 522 | |||||
| 557 | |||||
| 579 | |||||
| 606 | |||||
| 650 | |||||
| 678 | |||||
| 1028 | |||||
| 1089 | |||||
| 1193 | |||||
| 1199 | |||||
| mean |
@@
In Fig. 4 we observe a similar behavior for the real-world datasets with eggp covering an area close to or better than the best competing algorithm. The largest difference was on the friction dataset in which eggpmo covered an area smaller th Operon. Considering AUC, eggpmo, Operon, eggpso obtained a similar average score and were ranked in this order. PySR and tinyGP obtained significatively worse average scores. Considering the ranks on the median of the (Table 4), the multi-objective version consistently achieved second place, but in this set, Operon was ranked first for most of the datasets. We should notice, though, that eggp results were always close to the best competing algorithm, while even Operon misbehaved in two datasets (flow and niku-2). When observing the statistical test results, we can conclude that there are no significant differences between the eggpmo and Operon, but the hypotheses of eggp being equivalent to PySR and tinyGP can be rejected. Regarding the model size, for this set of benchmarks in which we used a smaller maximum size, all algorithms returned a model size close to the maximum, on average. In this criteria, PySR consistently returned smaller models but with the expense of smaller accuracy as seen in the AUC results.
| dataset | eggpmo | eggpso | Operon | PySR | tinyGP |
| chemical 1 | |||||
| chemical 2 | |||||
| flow | |||||
| friction | |||||
| friction dyn | |||||
| nasa 1 | |||||
| nasa 2 | |||||
| niku 1 | |||||
| niku 2 | |||||
| mean | |||||
| -value | 0.02 | 0.93 | 0.00 | 0.01 | |
| chemical 1 | |||||
| chemical 2 | |||||
| flow | |||||
| friction | |||||
| friction dyn | |||||
| nasa 1 | |||||
| nasa 2 | |||||
| niku 1 | |||||
| niku 2 | |||||
| mean | |||||
| -value | 0.00 | 0.85 | 0.00 | 0.00 | |
| chemical 1 | |||||
| chemical 2 | |||||
| flow | |||||
| friction | |||||
| friction dyn | |||||
| nasa 1 | |||||
| nasa 2 | |||||
| niku 1 | |||||
| niku 2 | |||||
| mean | |||||
| mean |
@@
Regarding the computational runtime, since Operon is the fastest symbolic regression implementation, as noted in (Burlacu et al., 2020b), we calculated the ratio between the average runtime of each algorithm to Operon. Fig. 5 shows the relative runtime per dataset. In this plot we can see that eggp and tinyGP were both between to times slower than Operon. PySR varied from to times the runtime of Operon depending on the dataset. The higher ratios were observed on high-dimensional or larger datasets. We should stress that all algorithms were constrained to run with a single thread, thus both Operon and PySR runtime could be smaller when exploiting multi-threading.
6.1. Additional considerations
The main limitations of these experiments lie in the use of default or reasonable values for the hyperparameters. We should notice that eggp contains hyperparameters that should be fine-tuned to obtain optimal results in a practical scenario. Meanwhile, PySR contains about hyperparameters that may affect the algorithm performance444we are not considering hyperparameters that can be set using prior knowledge. and Operon contains about hyperparameters. With careful experimentation, both PySR and Operon could have achieved similar results to those obtained with our approach. Having said that, eggp is a step forward to a parameterless experience in SR implementations, where the user only needs to set hyper-parameters that are intuitive w.r.t. their behavior. For example, increasing the number of evaluations will never make the results worse, unlike crossover and mutation rates which behaves unpredictably.
Another feature of eggp is the ability to export the current state of the e-graph into a file and load it when starting a new search. When loading a previous e-graph, eggp will resume the search by first extracting an initial population from that e-graph. In such case, eggpmo will recover the current Pareto front and eggpso will sample random solutions. The search can be resumed with different hyper-parameters, so the user can include new non-terminals or increase the maximum size. The e-graph file can also be used with the exploration tool rEGGression(de Franca and Kronberger, 2025) that allows the user to explore the history of solution and retrieve a regression model outside of the Pareto front.
6.2. Data availability
The algorithm is implemented in Haskell using the srtree555https://github.com/folivetti/srtree library for symbolic regression and equality saturation. The binaries and source code of eggp and tinyGP used in this paper are available at https://github.com/folivetti/srtree/releases/tag/v2.0.1.0 and all the datasets, scripts, and results to replicate the experiments are available at https://github.com/folivetti/eggp_paper_GECCO.
@@ 7. Conclusions
In this paper we explored the use of e-graphs and equality saturation as a mechanism to keep track of the history of the search engine for symbolic regression and to exploit its pattern matching capabilities to propose a variation to the traditional subtree crossover and mutation that increases the probability of generating novel expressions.
The e-graph data structure compactly stores shared elements of a set of expressions and their equivalence relationships, and efficiently queries for parts of expressions. Exploiting this capability, we modified the subtree operators to only sample subtrees that would generate unvisited expressions. The expectation is that this simple modification would render a significant improvement in the search procedure.
We tested the proposed algorithm, called eggp, in different benchmarks from the literature and compared with a simple GP using the original subtree operators, and two state-of-the-art algorithms, PySR and Operon. The results showed that the modified operators are capable of improving the performance of a simple GP to compete with the state-of-the-art. The main highlight of this approach is that it consistently performs equal or better than the best competing approach.
Regarding the runtime, eggp is consistently faster than PySR but significantly slower than Operon. When comparing with tinyGP, we can see that the use of e-graph and equality saturation does not increase the runtime significantly.
In conclusion, the expressiveness and capabilities of the e-graph data structure enabled us to make a simple modification to the original subtree operators while significantly improving the performance of GP for symbolic regression. This allowed us to obtain a more robust algorithm delivering better performance more reliably than state-of-the-art implementations. As for the next steps, this same modifications can be applied to any other operator used by the state-of-the-art algorithms. Not only that, but the e-graph opens up many new possibilities for improving the search as it allows us to query expressions with a combination of properties, which can translate to diversity-preserving population and easy to integrate prior knowledge (Kronberger et al., 2022; Reuter et al., 2024). In addition, the storage of the search history allows us to analyze the learned building blocks and exploit this information to generate new solutions.
@@ Acknowledgements.
F.O.F. is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) grant 301596/2022-0. G.K. is supported by the Austrian Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation and Technology, the Federal Ministry for Labour and Economy, and the regional government of Upper Austria within the COMET project ProMetHeus (904919) supported by the Austrian Research Promotion Agency (FFG). The authors proudly made no use of LLMs for this work.References
@@
- Bartlett et al. (2024) Deaglan J. Bartlett, Harry Desmond, and Pedro G. Ferreira. 2024. Exhaustive Symbolic Regression. IEEE Transactions on Evolutionary Computation 28, 4 (2024), 950–964. https://doi.org/10.1109/TEVC.2023.3280250 arXiv:2211.11461 [astro-ph.CO]
- Burlacu et al. (2019) Bogdan Burlacu, Michael Affenzeller, Gabriel Kronberger, and Michael Kommenda. 2019. Online Diversity Control in Symbolic Regression via a Fast Hash-based Tree Similarity Measure. In 2019 IEEE Congress on Evolutionary Computation (CEC). 2175–2182. https://doi.org/10.1109/CEC.2019.8790162
- Burlacu et al. (2020a) Bogdan Burlacu, Lukas Kammerer, Michael Affenzeller, and Gabriel Kronberger. 2020a. Hash-Based Tree Similarity and Simplification in Genetic Programming for Symbolic Regression. In Computer Aided Systems Theory – EUROCAST 2019, Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia (Eds.). Springer International Publishing, Cham, 361–369.
- Burlacu et al. (2020b) Bogdan Burlacu, Gabriel Kronberger, and Michael Kommenda. 2020b. Operon C++: an efficient genetic programming framework for symbolic regression. In Proceedings of the 2020 Genetic and Evolutionary Computation Conference Companion (Cancún, Mexico) (GECCO ’20). Association for Computing Machinery, New York, NY, USA, 1562–1570. https://doi.org/10.1145/3377929.3398099
- Cao et al. (2024) Haotian Cao, Garrett W Merz, Kyle Cranmer, and Gary Shiu. 2024. Learning Conformal Field Theory with Symbolic Regression: Recovering the Symbolic Expressions for the Energy Spectrum. In NeurIPS 2024 Workshop: Machine Learning and the Physical Sciences.
@@
- de Franca and Kronberger (2025) Fabricio Olivetti de Franca and Gabriel Kronberger. 2025. rEGGression: an Interactive and Agnostic Tool for the Exploration of Symbolic Regression Models. In Proceedings of the Genetic and Evolutionary Computation Conference (Malaga, Spain) (GECCO ’25). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3712256.3726385 arXiv:2501.17859 [cs.LG]
- de Franca et al. (2024) F. O. de Franca, M. Virgolin, M. Kommenda, M. S. Majumder, M. Cranmer, G. Espada, L. Ingelse, A. Fonseca, M. Landajuela, B. Petersen, R. Glatt, N. Mundhenk, C. S. Lee, J. D. Hochhalter, D. L. Randall, P. Kamienny, H. Zhang, G. Dick, A. Simon, B. Burlacu, Jaan Kasak, Meera Machado, Casper Wilstrup, and W. G. La Cavaz. 2024. SRBench++: Principled Benchmarking of Symbolic Regression With Domain-Expert Interpretation. IEEE Transactions on Evolutionary Computation (2024), 1–1. https://doi.org/10.1109/TEVC.2024.3423681
- Deb et al. (2000) Kalyanmoy Deb, Samir Agrawal, Amrit Pratap, and T. Meyarivan. 2000. A Fast Elitist Non-dominated Sorting Genetic Algorithm for Multi-objective Optimization: NSGA-II. In Parallel Problem Solving from Nature PPSN VI, Marc Schoenauer, Kalyanmoy Deb, Günther Rudolph, Xin Yao, Evelyne Lutton, Juan Julian Merelo, and Hans-Paul Schwefel (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 849–858.
- Ebner et al. (2001) Marc Ebner, Mark Shackleton, and Rob Shipman. 2001. How neutral networks influence evolvability. Complexity 7, 2 (2001), 19–33. https://doi.org/10.1002/cplx.10021 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1002/cplx.10021
- Gustafson et al. (2005) S. Gustafson, E.K. Burke, and N. Krasnogor. 2005. On improving genetic programming for symbolic regression. In 2005 IEEE Congress on Evolutionary Computation, Vol. 1. 912–919 Vol.1. https://doi.org/10.1109/CEC.2005.1554780
- Hu and Banzhaf (2018) Ting Hu and Wolfgang Banzhaf. 2018. Neutrality, Robustness, and Evolvability in Genetic Programming. Springer International Publishing, Cham, 101–117. https://doi.org/10.1007/978-3-319-97088-2_7
- Hu et al. (2023) Ting Hu, Gabriela Ochoa, and Wolfgang Banzhaf. 2023. Phenotype Search Trajectory Networks for Linear Genetic Programming. In Genetic Programming, Gisele Pappa, Mario Giacobini, and Zdenek Vasicek (Eds.). Springer Nature Switzerland, Cham, 52–67.
- Imai Aldeia et al. (2024) Guilherme Seidyo Imai Aldeia, Fabrício Olivetti De França, and William G. La Cava. 2024. Inexact Simplification of Symbolic Regression Expressions with Locality-sensitive Hashing. In Proceedings of the Genetic and Evolutionary Computation Conference (Melbourne, VIC, Australia) (GECCO ’24). Association for Computing Machinery, New York, NY, USA, 896–904. https://doi.org/10.1145/3638529.3654147
- Keller and Banzhaf (1999) Robert E. Keller and Wolfgang Banzhaf. 1999. The evolution of genetic code in Genetic Programming. In Proceedings of the 1st Annual Conference on Genetic and Evolutionary Computation - Volume 2 (Orlando, Florida) (GECCO’99). Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 1077–1082.
@@
- Kronberger et al. (2022) G. Kronberger, F. O. de Franca, B. Burlacu, C. Haider, and M. Kommenda. 2022. Shape-Constrained Symbolic Regression—Improving Extrapolation with Prior Knowledge. Evolutionary Computation 30, 1 (03 2022), 75–98. https://doi.org/10.1162/evco_a_00294
@@
- Kronberger and Olivetti de França (2025) Gabriel Kronberger and Fabrício Olivetti de França. 2025. Effects of reducing redundant parameters in parameter optimization for symbolic regression using genetic programming. Journal of Symbolic Computation 129 (2025), 102413. https://doi.org/10.1016/j.jsc.2024.102413
@@
- Moraglio et al. (2012) Alberto Moraglio, Krzysztof Krawiec, and Colin G Johnson. 2012. Geometric semantic genetic programming. In International Conference on Parallel Problem Solving from Nature. Springer, 21–31.
@@
- Ruberto et al. (2019) Stefano Ruberto, Leonardo Vanneschi, and Mauro Castelli. 2019. Genetic programming with semantic equivalence classes. Swarm and evolutionary computation 44 (2019), 453–469.
@@
- Schmidt and Lipson (2009) Michael Schmidt and Hod Lipson. 2009. Distilling free-form natural laws from experimental data. science 324, 5923 (2009), 81–85.
@@
- Uy et al. (2010) Nguyen Quang Uy, Michael O’Neill, Nguyen Xuan Hoai, Bob Mckay, and Edgar Galván-López. 2010. Semantic similarity based crossover in GP: The case for real-valued function regression. In Artifical Evolution: 9th International Conference, Evolution Artificielle, EA, 2009, Strasbourg, France, October 26-28, 2009. Revised Selected Papers 9. Springer, 170–181.
- Vanneschi et al. (2014) Leonardo Vanneschi, Mauro Castelli, and Sara Silva. 2014. A survey of semantic methods in genetic programming. Genetic Programming and Evolvable Machines 15 (2014), 195–214.
- Virgolin and Pissis (2022) Marco Virgolin and Solon P Pissis. 2022. Symbolic Regression is NP-hard. Transactions on Machine Learning Research (2022). https://openreview.net/forum?id=LTiaPxqe2e